OpenStudy (anonymous):

Ginny's class ordered 3 pizzas. They ate 5/12 of each pizza .How many pizzas did they eat?

OpenStudy (anonymous):

If they 5/12 of each pizza, then they didn't eat atleast one complete pizza. This is a logical answer Now, the mathematical answer is: they ate 5/12 of each pizza and there are 3 pizzas. So, they ate 5/12 x 3 = 5/4 = 1.25 pizza